The Portland Division's operating data for the past two years is as follows:

  Year 1  Year 2

Return on investment 12% 33%

Net operating income$? $495,000

Turnover ?  2

Margin ?  ?

Sales$2,320,000  ?


The Portland Division's margin in Year 2 was 150% of the margin for Year 1.

The turnover for Year 1 was:

Noreen 4e Rechecks 2017-31-03, 2017-17-04

rev: 07_17_2019_QC_CS-173103

13.00

1.96

1.09

3.00

Last year the Uptown Division of Gorcen Enterprises had sales of $645,000 and a net operating income of $70,950. The average operating assets at Uptown last year amounted to $284,000.

Last year at Uptown the margin used to calculate ROI amounted to:

11%

15%

25%

44%
Answer #1
1. calculate the turnover for Year 1 as follows:
Year 1 Year 2
Return on investment 12% 33%
Net operating income 255200 495000
Average operating assets 2126667 1500000
Turnover 1.09 2
Margin 11.0% 16.5%
Sales 2320000 3000000
Therefore,
The turnover for Year 1 was 1.09
2. Calculate the margin used to calculate ROI last year at Uptown as follows:
Net operating income 70950
Sales 645000
Margin 11%
Therefore,
Last year at Uptown the margin used to calculate ROI amounted to 11%

Above figures have been calculated in the following manner:

1. calculate the turnover for Year 1 as follows:
Year 1 Year 2
Return on investment 0.12 0.33
Net operating income =2320000*11% 495000
Average operating assets =255200/12% =495000/33%
Turnover =2320000/2126667 2
Margin =16.5%/150% =495000/3000000
Sales 2320000 =1500000*2
Therefore,
The turnover for Year 1 was 1.09
2. Calculate the margin used to calculate ROI last year at Uptown as follows:
Net operating income 70950
Sales 645000
Margin =B18/B19
Therefore,
Last year at Uptown the margin used to calculate ROI amounted to 11%